OpenStudy (anonymous):

9=|x+5| please help, having truble

jimthompson5910 (jim_thompson5910):

Hint: if |x| = k, then x = k or x = -k where k is some positive number So in your case, 9=|x+5| can be written as |x+5|=9 and that breaks down into the two equations x+5 = 9 OR x+5 = -9 Solve these last two equations to find your answers